# thanks: Denotes Equal Contributionthanks: Denotes Equal Contribution

### v2

@@ 2.1.1 Bootstrap Embedding
Utilizing the Schmidt Decomposition, we then construct the entangled bath orbitals for the given fragment. We build an exactly reduced wavefunction around fragment $F$, with fragment orbitals |$\ket{f^{F}}$ and entangled bath orbitals |$\ket{b^{F}}$. This procedure also results in unentangled environment orbitals |$\ket{e^{F}}$, which span the rest of the space. This reduced wavefunction in Equation 2 has a dimension of at most $2N_{\text{frag}}^{F}$ no matter the size of the full system.
@@
We perform correlated calculations using high-level wavefunction methods such as CCSD, FCI, and DMRG within the formulated fragment space after constructing the interacting bath Hamiltonian.[50] For self-consistency across multiple fragments, we optimize the global chemical potential which ensures a correct number of electrons in the chemical system, as discussed in previous works.[53, 54] We construct the energy, using a cumulant-based definition introduced in previous publication[66], for the system by taking components of the reduced density matrix, cumulant, and Hamiltonian from each fragment calculation, careful to avoid overcounting.[68, 66, 69]
@@ 2.1.2 Bootstrap Embedding with MM Charges
$\hat{H}_{\text{QMMM}}$ $=\sum_{\mu\nu}^{N_{\text{basis}}}(h_{\mu\nu}+U_{\mu\nu})\hat{c}_{\mu}^{\dagger}\hat{c}_{\nu}+\frac{1}{2}\sum_{\mu\nu\lambda\sigma}^{N_{\text{basis}}}V_{\mu\nu\lambda\sigma}\hat{c}_{\mu}^{\dagger}\hat{c}_{\lambda}^{\dagger}\hat{c}_{\sigma}\hat{c}_{\nu}$ (4)
@@ 3.1 Automatic QM Region Selection
When running large-scale high-throughput QM/MM or QM/QM/MM calculations for ensembles of large biological systems, manual division between the QM and MM regions becomes cumbersome. Here, we present an automatic procedure that identifies QM regions for a set of structures sampled from an MD trajectory. To generate a QM region, we introduce a variant of the automated QM region construction algorithm proposed in Reference 72. The original approach [72] generates a QM region such that the forces calculated with a QM/MM model for a selected atom matches the forces calculated for a very large reference QM region with the constraint of a maximum number of atoms $n^{\mathrm{region}}_{\mathrm{max}}$ in the final QM region. In our updated version of the algorithm, we perform this analysis for all ligand atoms, rather than for a single atom. Furthermore, we consider not only a single molecular structure to generate the QM region but a set of structures sampling the accessible configuration space at given conditions. This procedure results in a large QM region, surrounded by MM point charges, describing the biological system of interest at multiple points in the trajectory. We describe this procedure in detail in the Supporting Information.
@@ 3.2 Biological Target System Preparation
We calculate the forces necessary for the automatic QM region selection procedure with the semi-empirical tight binding method GFN2-xTB [75]. As discussed in Section 3.1, we extract a total of 100 snapshots equally spaced in time from the MM trajectory of the MCL1-19G protein–ligand complex of Reference 74. The parameters discussed in the Supporting Information for the QM region selection are chosen as follows: We set the maximum size of the reference QM region $n^{\mathrm{ref}}_{\mathrm{max}}$ to $500$ atoms, the maximum size of the candidate QM regions to $n^{\mathrm{region}}_{\mathrm{max}}=250$ atoms, and the target QM region size to $n^{\mathrm{region}}_{\mathrm{target}}=300$ atoms. Furthermore, we choose the minimum value for the search radius $r_{s}$ as $4\,$\mathrm{a}\mathrm{u}$$, the cutting probability $p_{\mathrm{cut}}$ as $0.9$, and the number of reference QM regions $N^{\mathrm{ref}}$ as $4$. The radius $r_{s}$ is incremented in steps of $0.1\,$\mathrm{a}\mathrm{u}$$, and $50$ candidate QM regions are generated for each value of $r_{s}$.
To avoid unphysical fragments in the QM calculations, we cap the QM region with hydrogen atoms at all bonds where a covalent bond was cut between the QM and MM regions. We account for the polarization of the QM region by the MM region through electrostatic embedding[39] of the QM region by the MM point charges. To avoid charge leaking from the QM region into the MM region caused by closely positioned MM charges, the charges are redistributed at the hydrogen caps based on the $RC$ scheme in Reference 76. We use the program SWOOSE[77, 78, 79] to perform the capping of the QM region and redistribute the MM point charges. We take the MM point charges from the Amber99sb*ILDN protein force field[80] and the TIP3P water model[81] employed in the original MM-MD simulation[74].
@@ 3.3 BE Calculations
We perform all BE calculations using QuEmb[71, 82], a Python package based on PySCF[70]. Github repository mixed-basis-be[83] includes both QM/QM/MM routines and mixed-basis BE functions such that the calculations introduced in this work can be replicated. For all calculations, we use CCSD as the high-level solver, though other wavefunction methods can be used.
@@ 4.2 Mixed-Basis BE for Polyacetylene and Polyglycine Chains
Figure 5 shows the percent error in the correlation energy compared to the standard BE results. That is, for BE$(n)$-in-BE$(m)$, we evaluate the percent error as follows, where CE is the correlation energy.
We observe that there is a robust convergence of BE$(1)$-in-BE$(m)$ with increasing $m$ (large basis set region size) to the standard BE$(1)$ result, despite the reduced computational cost from using a small fraction of the basis functions. BE$(2)$-in-BE$(2)$ already closely approximates the BE$(2)$ correlation energies, reaching about 1.5% error from BE$(2)$. For the tested systems, mixed-basis calculations provide a robust approximation to the sub-1% errors of large basis BE$(2)$ calculations relative to CCSD references. We show in Figure 6 that the convergence behavior remains comparable for the different choices of the basis set and the chain length. We note that BE$(n)$-in-BE$(n)$ has a notable degradation with larger basis and longer chain length. We discuss in Section 5.2 this behavior specifically in the case of BE$(1)$-in-BE$(1)$ and propose future work to reduce the deviation.
@@ 4.3 Mixed-Basis BE for Large Biological Systems
While we present here the single-point correlation energies of an ensemble of structures, the energies obtained with this QM/QM/MM approach can be further used alongside non-equilibrium switching[84] and other techniques to obtain the system binding free energies. To this end, this correlation and its more accurate calculation are crucial to understand drug binding, where interactions beyond the mean-field can be important. Per our benchmark studies, the estimation of these energies improves with increasing fragment size as quantum hardware improves, leading to a more accurate depiction of systems of interest with increased computational resources.
@@ 5.1 Mixed-Basis BE Scaling
Figure 9: Empirical scaling for $C_{16}H_{18}$ HF calculations and integral rotations. This is the total time for all reference HF calculations and for all integral transformations into all fragment spaces for the given system. All timings are fit to the form $t=A*r^{4}+B*r^{3}+C$, except for the BE$(n)$ integral transformation timings fit to $t=A*r^{B}+C$.
@@
As a demonstrative example, we will focus on the timing results for a simple, homogeneous, 16 carbon E-polyacetylene chain. This allows for us to obtain a complete set of reference BE$(n)$ timings with relatively large basis sets, which are compared with BE$(n)$-in-BE$(m)$ timings in various basis sets in Figure 9. For both the HF calculation and the formation of the fragment Hamiltonians, the bottleneck is four-index integral transformations, whose computational cost formally scales as $\mathcal{O}(N_{\text{bf}}^{4})$. Accordingly, if one moves from a STO-3G basis ($N_{\text{bf}}\approx 4$) to a cc-pVTZ basis ($N_{\text{bf}}\approx 30$) for standard large-basis BE$(n)$, one expects a calculation of both the HF energy and the fragment Hamiltonians to slow by a factor of $\left(\frac{30}{4}\right)^{4}\approx 3000$, a substantial slowdown. Indeed, we observe a slowdown in both steps in the calculation as shown by the solid curves in Figure 9.
@@ 5.2 Electron Count and Self-consistency
Figure 10: Correlation plot of the percent errors in electron count and in CEs (relative to BE$(n)$) from mixed-basis BE calculations on polyacetylene chains

### v3

@@ 1 Introduction
We note that the successful application of quantum computers for computational chemistry requires various considerations beyond tunable control over the number of qubits. In particular, for workflows that utilize QPE as the eigensolver, state preparation is a key concern, as a guiding state with enough overlap must be generated and provided. To this end, previous work has shown that mean-field ground states serve as robust guiding states for BE Hamiltonians,[64] which reduces the required quantum circuit depth for robust phase estimation.[65, 66]
Still, to obtain sensible results with QPE, we require quantum devices which satisfy qubit size and circuit depth requirements. A related work on resource estimation shows that chemical Hamiltonians of moderate size at the upper end of large FCI calculations ($\sim$ 16 spatial orbitals) require about $10^{6}-10^{7}$ two-qubit gates.[67] The study predicts, based on device roadmaps, that a gate count of about $10^{5}$ would be attainable within the next 3 years. The work claims that the gap is modest enough with advances in error correction to anticipate QPE-based workflow for quantum embedding Hamiltonians like BE in the near future.[67] In this work, we propose strategies that mainly focus on two aspects of quantum computing for chemical applications: (i) number of available qubits and (ii) the interplay between classical high-performance computing and quantum computers.
@@ 2.1.1 Bootstrap Embedding
Utilizing the Schmidt Decomposition, we then construct the entangled bath orbitals for the given fragment. We build an exactly reduced wavefunction around fragment $F$, with fragment orbitals |$\ket{f^{F}}$ and entangled bath orbitals |$\ket{b^{F}}$. This procedure also results in unentangled environment orbitals |$\ket{e^{F}}$, which span the rest of the space. This reduced wavefunction in Equation 2 has a dimension of at most $2N_{\text{frag}}^{F}$, where $N_{\text{frag}}^{F}$ is the number of fragment orbitals in the fragment $F$, no matter the size of the full system. $\Lambda^{F}_{k}$ are the Schmidt coefficients associated with each state.
@@
We perform correlated calculations using high-level wavefunction methods such as CCSD, FCI, and DMRG within the formulated fragment space after constructing the interacting bath Hamiltonian.[50] For self-consistency across multiple fragments, we optimize the global chemical potential which ensures a correct number of electrons in the chemical system, as discussed in previous works.[53, 54] We construct the energy, using a cumulant-based definition introduced in previous publication[70] and in more detail in the Supporting Information, for the system by taking components of the reduced density matrix, cumulant, and Hamiltonian from each fragment calculation, careful to avoid overcounting.[72, 70, 73]
@@ 2.1.2 Bootstrap Embedding with MM Charges
$\hat{H}_{\text{QMMM}}=\sum_{\mu\nu}^{N_{\text{basis}}}(h_{\mu\nu}+U_{\mu\nu})\hat{c}_{\mu}^{\dagger}\hat{c}_{\nu}\\ +\frac{1}{2}\sum_{\mu\nu\lambda\sigma}^{N_{\text{basis}}}V_{\mu\nu\lambda\sigma}\hat{c}_{\mu}^{\dagger}\hat{c}_{\lambda}^{\dagger}\hat{c}_{\sigma}\hat{c}_{\nu}$ (4)
@@ 2.2 Mixed-Basis BE
It is important to note that mixed-basis approaches have been explored in depth atop other correlated methods over many years.[76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87] In this work, we use the BE systematic black-box framework alongside the mixed-basis set idea to clearly define each basis set region. Unlike many other methods, which choose some subset of orbitals to be in a denser basis set than the others, we describe the entire system with the chosen large basis set across the set of calculations. This approximates and indeed converges to the absolute energy of the full system in the chosen large basis set.
@@ 3.1 Automatic QM Region Selection
When running large-scale high-throughput QM/MM or QM/QM/MM calculations for ensembles of large biological systems, manual division between the QM and MM regions becomes cumbersome. Here, we present an automatic procedure that identifies QM regions for a set of structures sampled from an MD trajectory. To generate a QM region, we introduce a variant of the automated QM region construction algorithm proposed in Reference 88. The original approach [88] generates a QM region such that the forces calculated with a QM/MM model for a selected atom matches the forces calculated for a very large reference QM region with the constraint of a maximum number of atoms $n^{\mathrm{region}}_{\mathrm{max}}$ in the final QM region. In our updated version of the algorithm, we perform this analysis for all ligand atoms, rather than for a single atom. Furthermore, we consider not only a single molecular structure to generate the QM region but a set of structures sampling the accessible configuration space at given conditions. This procedure results in a large QM region, surrounded by MM point charges, describing the biological system of interest at multiple points in the trajectory. We describe this procedure in detail in the Supporting Information.
@@ 3.2 Biological Target System Preparation
We calculate the forces necessary for the automatic QM region selection procedure with the semi-empirical tight binding method GFN2-xTB [91]. As discussed in Section 3.1, we extract a total of 100 snapshots equally spaced in time from the MM trajectory of the MCL1-19G protein–ligand complex of Reference 90. The parameters discussed in the Supporting Information for the QM region selection are chosen as follows: We set the maximum size of the reference QM region $n^{\mathrm{ref}}_{\mathrm{max}}$ to $500$ atoms, the maximum size of the candidate QM regions to $n^{\mathrm{region}}_{\mathrm{max}}=250$ atoms, and the target QM region size to $n^{\mathrm{region}}_{\mathrm{target}}=300$ atoms. Furthermore, we choose the minimum value for the search radius $r_{s}$ as $4\,$\mathrm{a}\mathrm{u}$$, the cutting probability $p_{\mathrm{cut}}$ as $0.9$, and the number of reference QM regions $N^{\mathrm{ref}}$ as $4$. The radius $r_{s}$ is incremented in steps of $0.1\,$\mathrm{a}\mathrm{u}$$, and $50$ candidate QM regions are generated for each value of $r_{s}$. Often large QM regions are required to converge relative energies in QM/MM.[92, 93, 94] Since we will focus on the correlation treatment in the QM region rather than predicting observables, such as binding free energies, we selected the key parameters $n^{\mathrm{region}}_{\mathrm{max}}$ and $n^{\mathrm{ref}}_{\mathrm{max}}$ as a compromise between computational cost and accuracy. A dedicated uncertainty analysis will be required to propagate uncertainties from the QM region selection to the target properties (e.g., binding free energies).
To avoid unphysical fragments in the QM calculations, we cap the QM region with hydrogen atoms at all bonds where a covalent bond was cut between the QM and MM regions. We account for the polarization of the QM region by the MM region through electrostatic embedding[39] of the QM region by the MM point charges. To avoid charge leaking from the QM region into the MM region caused by closely positioned MM charges, the charges are redistributed at the hydrogen caps based on the $RC$ scheme in Reference 95. We use the program SWOOSE[96, 97, 88] to perform the capping of the QM region and redistribute the MM point charges. We take the MM point charges from the Amber99sb*ILDN protein force field[98] and the TIP3P water model[99] employed in the original MM-MD simulation[90].
@@ 3.3 BE Calculations
We perform all BE calculations using QuEmb[75, 100], a Python package based on PySCF[74]. Github repository mixed-basis-be[101] includes both QM/QM/MM routines and mixed-basis BE functions such that the calculations introduced in this work can be replicated. For all calculations, we use CCSD as the high-level solver, though other wavefunction methods in principle can be used.
@@ 4.2 Mixed-Basis BE for Polyacetylene and Polyglycine Chains
Figure 5 shows the percent error in the correlation energy compared to the standard BE results for molecules of increasing length with the cc-pVDZ basis set. That is, for BE$(n)$-in-BE$(m)$, we evaluate the percent error as follows, where CE is the correlation energy.
We observe that there is a robust convergence of BE$(1)$-in-BE$(m)$ with increasing $m$ (large basis set region size) to the standard BE$(1)$ result, despite the reduced computational cost from using a small fraction of the basis functions. BE$(2)$-in-BE$(2)$ already closely approximates the BE$(2)$ correlation energies, reaching about 1.5% error from BE$(2)$. For the tested systems, mixed-basis calculations provide a robust approximation to the sub-1% errors of large basis BE$(2)$ calculations relative to CCSD references. We provide geometry-averaged, per-carbon results in Figure 6 to show that the convergence behavior remains comparable for the different choices of the basis set. By showing per-carbon error, we can compare systems of different lengths across these basis sets, rather than showing a single calculation of fixed length at each, to evaluate the method behavior. We note that BE$(n)$-in-BE$(n)$ has a notable degradation with larger basis (shown in Figure 6) and longer chain length(Figure 5), which is discussed further in the Supporting Information alongside the full results data. We discuss in Section 5.2 this behavior specifically in the case of BE$(1)$-in-BE$(1)$ and propose future work to reduce the deviation.
@@ 4.3 Mixed-Basis BE for Large Biological Systems
While we present here the single-point correlation energies of an ensemble of structures, the energies obtained with this QM/QM/MM approach can be further used alongside non-equilibrium switching[102] and other techniques to obtain the system binding free energies. The free binding energy for this system requires extensive calculations beyond the scope of this work, but such a method has been described at length in other studies.[90, 103] To this end, this correlation and its more accurate calculation are crucial to understand drug binding, where interactions beyond the mean-field can be important. Per our benchmark studies, the estimation of these energies improves with increasing fragment size as quantum hardware improves, leading to a more accurate depiction of systems of interest with increased computational resources.
@@ 5.1 Mixed-Basis BE Scaling
Figure 9: Empirical scaling for $C_{16}H_{18}$ HF calculations and integral rotations. This is the total time for all reference HF calculations and for all integral transformations into all fragment spaces for the given system. BE$(n)$ timings are fit with a solid curve, while BE$(n)$-in-BE$(m)$ are fit with dotted curves. BE$(1)$-in-BE$(1)$ is highlighted in orange, BE$(1)$-in-BE$(2)$ is highlighted purple, and BE$(1)$-in-BE$(3)$ is highlighted green. For integral transformation timings, BE$(m)$ and BE$(n)$-in-BE$(m)$ data share a marker shape. All timings are fit to the form $t=A*r^{4}+B*r^{3}+C$, except for the BE$(n)$ integral transformation timings fit to $t=A*r^{B}+C$.
@@
As a demonstrative example, we will focus on the timing results for a simple, homogeneous, 16 carbon E-polyacetylene chain. This allows us to obtain a complete set of reference BE$(n)$ timings with relatively large basis sets, which are compared with BE$(n)$-in-BE$(m)$ timings in various basis sets in Figure 9. For both the HF calculation and the formation of the fragment Hamiltonians, the bottleneck is four-index integral transformations, whose computational cost formally scales as $\mathcal{O}(N_{\text{bf}}^{4})$. Accordingly, if one moves from a STO-3G basis ($N_{\text{bf}}\approx 4$) to a cc-pVTZ basis ($N_{\text{bf}}\approx 30$) for standard large-basis BE$(n)$, one expects a calculation of both the HF energy and the fragment Hamiltonians to slow by a factor of $\left(\frac{30}{4}\right)^{4}\approx 3000$, a substantial slowdown. Indeed, we observe a slowdown in both steps with changing basis set in the calculation as shown by the solid curves, representing BE$(n)$ data, in Figure 9. Note that the integral transformation steps, in solid blue, are similarly expensive for BE$(1)$, BE$(2)$, and BE$(3)$ for a given basis set.
@@ 5.2 Electron Count and Self-consistency
Figure 10: Correlation plot of the percent errors in electron count and in CEs (relative to BE$(n)$) from mixed-basis BE calculations on polyacetylene chains. We see a strong linear relationship between electron count error and percent CE error for BE$(1)$-in-BE$(1)$ relative to BE$(1)$ and a much weaker relationship for other mixed-basis set calculations.